### v2

# DIRECT: Deep Active Learning under Imbalance and Label Noise

Abstract
Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes. With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation. Label noise is another common issue in data annotation jobs, which is especially challenging for active learning methods. In this work, we conduct the first study of active learning under both class imbalance and label noise. We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it. Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage the classic active learning literature to address issues such as batch labeling and tolerance towards label noise. We present extensive experiments on imbalanced datasets with and without label noise. Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
Machine Learning, ICML
1 Introduction
Large-scale deep learning models are playing increasingly important roles across many industries. Human feedback and annotations have played a significant role in developing such systems through. Progressively over time, we believe the role of humans in a machine learning pipeline would shift to annotating rare yet important cases. However, under data imbalance, the typical strategy of randomly choosing examples for annotation becomes especially inefficient. This is because the majority of the labeling budget would be spent on common and well-learned classes, resulting in insufficient rare class examples for training an effective model. For many annotation jobs, this challenge of data imbalance is further compounded by label noises induced from annotators. In this paper, we mitigate these issues by proposing a deep active learning strategy. Our algorithm DIRECT sequentially and adaptively chooses informative and more class-balanced examples for annotation while being robust to noisy annotations. To the best of our knowledge, this is the first deep active learning study to address the challenging yet prevalent scenario where both imbalance and label noise coexist.
Inspired by previous work by Zhang et al. (2022), we propose a novel reduction of the imbalanced classification problem into one-dimensional active learning problems. For each class, our reduction sorts unlabeled examples into an list ordered by one-vs-rest margin scores. The goal of DIRECT is to find the optimal separation threshold which best separates the examples in any given class from the rests. With an equivalence to finding the best threshold classifier, we were able to employ classic active learning literature for obtaining the separation threshold. Once the optimal separation threshold is identified, DIRECT annotates a set of more class-balanced examples near this threshold. As examples around the optimal separation threshold tend to have high uncertainty, DIRECT annotates examples that are both class-balanced and informative.
With the reduction to classic one-dimensional active learning literature, DIRECT also inherits the advantages from classic active learning algorithms such as robustness against label noise and batch labeling. This is in contrast to the state-of-art imbalanced active learning algorithm GALAXY proposed by Zhang et al. (2022), which requires sequential labeling and is much less robust against label noise. As demonstrated in Figure 1, under both noisy and noiseless label settings, DIRECT improves significantly over GALAXY’s label efficiency – DIRECT requires less annotations to achieve the same accuracy. It achieves state-of-art label efficiency gain on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench framework (Zhang et al., 2023a). Lastly, DIRECT also allows for parallel annotations with various batch sizes.
(a) Imbalanced CIFAR-10, two classes, no label noise.
(b) Imbalanced CIFAR-100, two classes, 20% label noise.
(c) LabelBench FMoW (62 classes with imbalance), no label noise
Figure 1: Performance of DIRECT over existing baselines for both noiseless and noisy settings. The x-axis represents the total number of labeled examples so far and the y-axis shows the neural network’s balanced accuracy. Both (a) and (b) are using supervised training of ResNet-18. In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2023a). $B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators for the annotation job. $B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY. Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world decisions.
To summarize our main contributions:
  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification for deep neural networks.
  • Our algorithm DIRECT addresses the prevalent imbalance and label noise issues and annotates a more class-balanced and informative set of examples.
  • Compared to state-of-art algorithm GALAXY (Zhang et al., 2022), DIRECT allows parallel annotation by multiple annotators while still maintaining significant label-efficiency improvement.
  • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32). DIRECT consistently performs the best compared against existing baseline algorithms. Quantitatively, DIRECT can save more than $60\%$ annotation cost compared to GALAXY, and more than $80\%$ annotation cost compared to random sampling.
2 Related Work
Class-Balanced Deep Active Learning Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009). Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have also been shown to perform among top methods when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2023a). More sophisticated methods have also been proposed to optimize the chosen unlabeled examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as under label noises (Khosla et al., 2022) or under class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2023a).
Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications. Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2023b). To this end, Kothawade et al. (2021) propose a submodular-based method that actively annotates examples similar to known examples of rare instances. Our work is motivated by Zhang et al. (2022), where they propose GALAXY, an algorithm that construct one-dimensional linear graphs and apply graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain. While GALAXY outperforms existing algorithms, due to a bisection procedure involved, it does not allow parallel annotation. In addition, bisection procedures are generally not robust against label noises, a prevalent challenge in real world annotation tasks. Our algorithm DIRECT mitigates all of the above shortcomings of GALAXY while outperforming it even with synchronous labeling and no label noise, beating GALAXY in its own game.
Lastly, we distinguish our work from Zhang et al. (2023b), where the paper studies the algorithm selection problem. Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature. Related but tangential to our work, several papers have studied to use active learning for cleaning existing noisy labels (Lin et al., 2016; Younesian et al., 2021). In this line of work, they assume access to an oracle annotator that will provide clean labels when queried upon. This is fundamentally different from our work, where our annotator may provide noisy labels. Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang & Chaudhuri, 2015; Chen et al., 2022). The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on those instances. In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today. Recently, Khosla et al. (2022) proposed a novel deep active learning algorithm specialized for Heteroskedastic noise, where different “regions” of examples are subject to different levels of noise. Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels.
To the best of our knowledge, no existing deep active learning work has studied the scenario where both imbalance and label noise present. Yet, this setting is the most prevalent in real-world annotation applications.
Agnostic Active Learning A critical part of our solution is the reduction to a one dimensional agnostic active learning problem, e.g. Section 4.1. There is extensive work on agnostic learning in general, and we refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion. All of these works, beginning with the seminal works (Balcan et al., 2006; Dasgupta et al., 2007) follow, a familiar paradigm of disagreement based learning. This involves maintainin a version space of promising hypothesis at each time, and constructing the disagreement region as a set of unlabeled examples for which there exists two hypotheses that disagree on their predictions. The example to label is chosen by sampling from a informative sampling distribution computed over the disagreement region. Several approaches have been proposed for computing the sampling distribution, e.g. Jain & Jamieson (2019); Katz-Samuels et al. (2020, 2021); Huang et al. (2015). As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al. (2021). We remark that the problem of class imbalance has been largely ignored by the classical work on traditional agnostic learning, and this paper hopes to close this gap.
3 Preliminary
3.1 Notations
We study the pool-based active learning problem, where an initial unlabeled set of $N$ examples $X=\{x_{1},...,x_{N}\}$ are available for annotation. Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown. Furthermore, we study the multi-class classification problem, where the space of labels $\mathcal{Y}:=[K]$ is consisted $K$ classes. Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ of each class. We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k^{\prime}\in[K]}N_{k^{\prime}}}$.
A deep active learning algorithm iteratively chooses batches of examples for annotation. During the $t$-th iteration, the algorithm is given labeled and unlabeled sets of examples, $L_{t}$ and $U_{t}$ respectively, where $L_{t}\cup U_{t}=X$ and $L_{t}\cap U_{t}=\emptyset$. The algorithm then chooses $B$ examples from the unlabeled set $X^{(t)}\subseteq U_{t}$ and then obtains their corresponding labels $Y^{(t)}$. The labeled and unlabeled sets are then updated, i.e., $L_{t+1}\leftarrow L_{t}\cup X^{(t)}$ and $U_{t+1}\leftarrow U_{t}\backslash X^{(t)}$. Based on new labeled set $L_{t+1}$ and its corresponding labels, a neural network $f_{t}:X\rightarrow[K]$ is trained to inform the choice for the next iteration. The ultimate goal of deep active learning is to obtain high predictive accuracy for the trained neural network while annotating as few examples as possible.
3.2 Review of Imbalanced Active Learning: GALAXY
We first consider a binary case of imbalanced classification, where we assume $N_{1}<N_{2}$ without loss of generality. After a total of $T$ rounds and sampling $TB$ examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$. On the other hand, as shown previously in Zhang et al. (2022), uncertainty based methods such as confidence (Settles, 2009), margin (Tong & Koller, 2001; Balcan et al., 2006) and entropy (Kremer et al., 2014) sampling could annotate a set of examples with imbalance ratio arbitrarily low. The reasoning is simply demonstrated in Figure 1(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples almost all from the majority class. It is also empirically shown that annotating examples most certain to be in the minority class is not the most effective in improving model performance.
[…]
(b) GALAXY spends approximately equal annotation budget around both cuts, while the cut on the right would yield examples mostly in the majority class.
Figure 2: Figures demonstrating GALAXY. Ordered lists of examples are ranked by the predictive sigmoid score $\widehat{p}$. The ground truth label of each example is represented by its border – solid blue lines for class $1$ and dotted red lines for class $2$. Annotated examples are shaded.
Therefore, the key objective in imbalanced active learning is to label examples that are both uncertain and class balanced. In both GALAXY (Zhang et al., 2022) and this work, we decompose the objective into the following two-phased procedure: 1. Identify the optimal separation threshold $j^{\star}$ that best separates the minority class from the majority class, as shown in Figure 1(a). 2. Annotate equal number of examples next to $j^{\star}$ from both sides.
GALAXY draws inspiration from graph-based active learning in identifying $p^{\star}$. In particular, it follows a modified bisection procedure that sequentially annotates examples to find all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 1(b)). As the optimal separation threshold must be located at one of these cuts, GALAXY annotates examples close to all the found cuts. However, the algorithm suffers from three weaknesses. First, as shown in Figure 1(b), by annotating around all cuts, the algorithm may waste a significant portion of the annotation budget around outliers, which could lead to annotating large number of majority class examples. Second, the above problem can be magnified when annotators provide noisy labels. Third, a bisection procedure only allows for sequential labeling, which prevents multiple annotators labeling in parallel. In this paper, we take a DIRECT approach by identifying only the optimal separation threshold and address all of the shortcomings above.
4 A Robust Algorithm with Batch Labeling
In this section, we formally define the optimal separation threshold and pose the problem of identifying it as an 1-dimensional reduction to the agnostic active learning problem. We then propose an algorithm inspired by the agnostic active learning literature (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021).
4.1 An 1-D Reduction to Agnostic Active Learning
We start by considering the imbalanced binary classification setting mentioned in Section 3.2. When given a neural network model, we let $\widehat{p}:X\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores. Here, a higher sigmoid score represents a higher confidence of the example being in class $2$. We sort examples by their sigmoid predictive score similar to Section 3.2. Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots\leq q_{(N)}$ be a sorted permutation of $\widehat{p}(x_{1}),...,\widehat{p}(x_{N})$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels. We are now ready to define the optimal separation threshold as described in Section 3.2.
Definition 4.1.
Given a sorting of $X$ the points according the sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
$j^{\star}=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=1:i\leq j\}\rvert-\lvert\{y_{(i)}=2:i\leq j\}\rvert\right)$
$=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=2:i>j\}\rvert-\lvert\{y_{(i)}=2:i>j\}\rvert\right).$ (1)
In other words, on either side of $j^{\ast}$, it has the largest discrepancy in the number of examples between the two classes. This captures the intuition of Figure 1(b) — our goal is to find a threshold that best separates one class from the other. We quickly remark that ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ is defined as
$h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}.$
and $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times. The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1). Namely, $j^{\star}\in\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$, transforming the problem of finding the optimal separation threshold to an 1-D agnostic active learning problem.
Multi-Class Classification. To generalize the above problem formulation to multi-class classification, we follow a similar strategy to Zhang et al. (2022). Specifically, for each class $k$, one can view the problem of class-$k$ v.s. others as a binary classification problem. In other words, class $k$ can be seen as class $1$ and classes $\{1,...,k-1,k+1,...,N\}$ can be seen as class $2$. The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems. Moreover, let $\widehat{p}:X\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores. Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\widehat{p}(x_{i})]_{k^{\prime}}-[\widehat{p}(x_{i})]_{k}$ to sort the examples. When multiple examples have the same margin score, we further break ties by their corresponding confidence scores $[\widehat{p}(x_{i})]_{k}$. We also note that sorting with margin scores is equivalent to sorting with sigmoid scores in the binary classification setting, so this is a strictly more general formulation to multi-class classification.
  Input: Pool $X$, number of rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
  Initialize: Uniformly sample $B$ elements without replacement from $X$ to form $L_{0}$. Let $U_{0}\leftarrow X\backslash L_{0}$.
  for $t=1,...,T-1$ do
     Train neural network on $L_{t-1}$ and obtain $f_{t-1}$.
     Find optimal separation thresholds
     Initialize labeled set $L_{t}\leftarrow L_{t-1}$.
     Initialize budget per class $b\leftarrow B_{\text{train}}/2K$.
     for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ and break ties so that $q_{(i)}^{k}=q_{(i+1)}^{k}\implies[\widehat{p}(x_{(i)})]_{k}\geq[\widehat{p}(x_{(i+1)})]_{k}$.
        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{q^{k}_{(i)}\}_{i=1}^{N})$.
     end for
     Annotate rare and uncertain examples
     Compute budget per class $b\leftarrow(B_{\text{train}}-|L_{t}|)/K$.
     for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
        Estimate separation threshold
$\widehat{j}^{k}\leftarrow$ $\operatorname*{arg\,max}_{j}(|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|$
|$-|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|)$
and break ties by choosing the index closest to $\frac{N}{2}$. Annotate $b$ unlabeled examples with sorted indices closest to $\widehat{j}^{k}$ and insert to $L_{t}$.
     end for
  end for
  Return: Train final classifier $f_{T}$ based on $L_{T}$.
Algorithm 1 DIRECT: DImension REduction for aCTive Learning under Imbalance and Label Noise
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{q_{(i)}\}_{i=1}^{N}$.
  Define: For examples in the labeled set, we let $y_{(i)}$ denote the label of the example corresponding to $q_{(i)}$.
  Initialize: Version space: shortest segment of indices $[I,J]$, with $q_{(I)},q_{(J)}$ $\forall i\leq I,q_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,q_{(j)}\in L\implies y_{(j)}\neq k$.
  Initialize: Number of iterations $m\leftarrow\frac{b}{B_{\text{parallel}}}$. Shrinking factor $c\leftarrow\sqrt[m]{J-I}$
  for $t=1,...,m$ do
     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $q_{(I)}$, …, $q_{(J)}$ for annotation and insert to $L$.
     Find $[I^{\prime},J^{\prime}]\subset[I,J]$ such that
$I^{\prime},J^{\prime}=\operatorname*{arg\,min}_{i,j:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}\quad\text{where }$
^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$
     Update version space $I,J\leftarrow I^{\prime},J^{\prime}$.
  end forReturn: Updated labeled set $L$.
Algorithm 2 VReduce: Version Space Reduction
4.2 Algorithm
We are now ready to state our algorithm as shown in Algorithm 1. The algorithm follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class. Our set of labeled examples at anytime are maintained in $L_{t}$.
To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2) on the set of labeled examples. Our method is motivated by the fixed-budget algorithm for active classification, ACED (Katz-Samuels et al., 2021). Intuitively, in the non-noisy or separable setting, where the loss of $\mathcal{L}(h_{j^{\star}})=0$, we could run a binary search procedure to learn $j^{\star}$. However, we can’t assume that $j^{\star}$ perfectly separates the classes. ACED circumvents this by maintaining a version space of threshold classifiers, namely an interval $[I,J]$. In each of the $b/B_{\text{parallel}}$ rounds, $B_{\text{parallel}}$ samples are taken and appended to the set of labeled examples $L$. Then, an interval $I^{\prime},J^{\prime}$ is computed, whose length is a factor of $1/c$ smaller than the current $I,J$ of low-loss classifiers. The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left. The fixed confidence (or alternatively $\delta$-PAC) analogue of ACED is known to match instance optimal lower bounds for active classification up to logarithmic factors. Under benign noise conditions, the probability of misidentifying the optimal separation threshold decays exponentially w.r.t. budget $b$.
To address batch labeling, we note the number of examples collected before retraining is usually far greater than the number of annotators annotating in parallel. In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained. In practice, this number is usually determined by the constraints of computational training cost. On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel. As discussed above, $B_{\text{parallel}}\ll B_{\text{train}}$ in practice. Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.
5 Experiments
We conduct experiments under two primary setups:
  1. 1.
    Fine-tuning ResNet-18 with vanilla supervised training on imbalanced datasets similar to Zhang et al. (2022).
  2. 2.
    Fine-tuning large pretrained model (CLIP ViT-B32) with semi-supervised training strategies under the LabelBench framework (Zhang et al., 2023a).
In Section 5.1, we document these setups in detail. For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.2, showing its superior label-efficiency and ability to accommodate batch labeling. In Section 5.3, we evaluate deep active learning algorithms under a novel setting with both class imbalance and noisy labels. Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises. While we highlight many results in this section, see Appendix B for all results.
5.1 Experiment Setup
ResNet-18 Experiments. ResNet-18 with passive training has been the standard evaluation in existing deep active literature (Ash et al., 2019; Zhang et al., 2022). Our experiment setup utilizes the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) image classification datasets. The original forms of these datasets are roughly balanced across 9, 10 or 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, given a balanced dataset above with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Imbalance ratios are shown in Table 1.
For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer (Kingma & Ba, 2014). The ResNet-18 model (He et al., 2016) is pretrained on ImageNet (Deng et al., 2009) from the PyTorch library. To address data imbalance, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class. For experiments with label noise, we further add a $10\%$ label smoothing during training (Müller et al., 2019).
LabelBench Experiments. Proposed by Zhang et al. (2023a), LabelBench evaluates active learning performance in a more comprehensive framework. Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model (Radford et al., 2021). The framework also utilizes semi-supervised learning method FlexMatch (Zhang et al., 2021) to further leverage the unlabeled examples in the pool for training. We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021). Similar to the ResNet-18 experiments, we use a $10\%$ label smoothing in the loss function to improve training under label noise. We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
Name # Classes $K$ $N$ Imb Ratio $\gamma=\frac{\min_{k}N_{k}}{\max_{k^{\prime}}N_{k^{\prime}}}$
Imb CIFAR-10 $2$ $50000$$1$.1111.1111.1111
Imb CIFAR-10 $3$ $50000$1250.1250.1250.1250
Imb CIFAR-100 $2$ $50000$0101.0101.0101.0101
Imb CIFAR-100 $3$ $50000$0102.0102.0102.0102
Imb CIFAR-100 $10$ $50000$0110.0110.0110.0110
Imb SVHN $2$ $73257$0724.0724.0724.0724
Imb SVHN $3$ $54448$2546.2546.2546.2546
PathMNIST $2$ $89996$1166.1166.1166.1166
FMoW $62$ $76863$0049.0049.0049.0049
iWildCam $14$ $129809$ $4.57\cdot 10^{-5}$
Table 1: Dataset settings for our experiments. $N$ denotes the total number of examples in our dataset. $\gamma$ is the class imbalance ratio defined in Section 3.1.
(a) Imbalanced CIFAR-10, three classes
(b) Imbalanced CIFAR-100, ten classes
(c) Imbalanced SVHN, two classes
(d) PathMNIST, two classes
(e) FMoW, Balanced Pool Accuracy
(f) Imbalanced SVHN, two classes, number of labeled examples in minority class
Figure 3: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting. (a)-(d) are balanced accuracy on ResNet-18 experiments while (e) shows experiment under the LabelBench framework. $B_{\text{parallel}}$ indicates the number of parallel annotations as mentioned in Section 4.2. $B_{\text{parallel}}=1$ is equivalent with sychornous labeling. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
5.2 Experiments under Imbalance, without Label Noise
For the noiseless experiment on ResNet-18, we compare against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. We briefly distinguish the algorithms into two categories. In particular, both SIMILAR and Most Likely Positive annotate examples that are similar to existing labeled minority examples, thus can significantly annotate a large quantity of minority examples. The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty. For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results. In particular, we compared against BADGE, Margin Sampling (Balcan et al., 2006; Kremer et al., 2014), CORESET and GALAXY.
As highlighted in Figures 1(a) and 3(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments. In Figures 1(c) and 3(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments. Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 3(c). In terms of class-balancedness, we consistently observe that both Most Likely Positive and SIMILAR annotating greater number of minority class examples, but significantly underperforms in terms of balanced accuracy (an example showin in Figure 3(f)). While Zhang et al. (2022) has already observed this phenomenon, we can further see that DIRECT collects slightly less minority class examples than GALAXY, but outperforms in terms of balanced accuracy. While it is crucial to optimize class-balancedness for better model performance, we see that both extremes of annotating too few and too many minority examples could lead to worse generalization performances. When too few examples are from minority class, the performance of the minority classes could be significantly hindered. When optimized to annotate as many examples from minority class as possible, the algorithm has to tradeoff annotating informative examples to examples it is more certain to be in the minority class. Together, this suggests an intricate balance between the two objectives, generalization performance and class-balancedness.
We would also like to highlight the ability to handle batch labeling. Across our experiments, we see DIRECT outperforms with different amounts of parallel annotation ($B_{\text{parallel}}=1,5$ and $20$), indicating its general effectiveness. This is in comparison to the synchoronous nature of GALAXY, where it is always using $B_{\text{parallel}}=1$. On Figures 1(a) and 4(a), we see that DIRECT outperforms GALAXY with synchronous labeling. Furthermore, in these experiments we also see using $B_{\text{parallel}}=5$ only affects algorithm performances minimally for DIRECT.
(a) Imbalanced CIFAR-10, two classes, 10% label noise
(b) Imbalanced SVHN, two classes, 10% label noise
(c) iWildcam Balanced Pool Accuracy, $10\%$ label noise
Figure 4: Performance of DIRECT against baseline algorithms under $10\%$ label noise. (a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
(a) Imbalanced CIFAR-100, two classes, no label noise
(b) Imbalanced CIFAR-100, two classes, 10% label noise
(c) Imbalanced CIFAR-100, two classes, 15% label noise
Figure 5: Performance of DIRECT against baseline algorithms under different levels of label noise. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
5.3 Experiments under Imbalance and Label Noise
We conduct novel sets of experiments under both class imbalance and label noise. Here, for both ResNet-18 and LabelBench experiments, we evaluate against all of the algorithms that performed well under the imbalance but noiseless setting above. For all of our experiments, we introduce a fixed percentage of label noise, where the given fraction of the examples’ labels are corrupted to a different class uniformly at random. For most of our experiements with $10\%$ label noise shown in Figure 4, we observe again that DIRECT consistently improves over all baselines including GALAXY. The results are consistent on ResNet-18 and LabelBench setups, and with different $B_{\text{parallel}}$ values, showing DIRECT’s robustness under label noise.
Different Levels of Label Noise As shown in Figures 5(a)-(c) and 1(b), we observe the results on imbalanced CIFAR-100 with two classes across numerous levels of label noise, with $0\%$, $10\%$, $15\%$ and $20\%$ respectively. In fact, the noiseless experiment in Figure 5(a) is the only setting DIRECT slightly underperforms GALAXY in terms of generalization accuracy. However, we see DIRECT becomes more label-efficient under the more practical label noise settings. It is also worth noting that with high label noise of $20\%$, we observe in Figure 1(b) that existing algorithms underperform random sampling. In contrast, DIRECT significantly outperforms random sampling, saving more than $60\%$ of the annotation cost.
6 Conclusion and Future Work
In this paper, we conducted the first study of deep active learning under both class imbalance and label noise. We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature. In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel. Studying asynchronous labeling could a natural extension of our work. We think a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm. In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators.
Acknowledgements
This work has been supported in part by NSF Award 2112471.
Impact Statement
In the rapidly evolving landscape of machine learning, the efficacy of active learning in addressing data imbalance and label noise is a significant stride towards more robust and equitable AI systems. This research explores how active learning can effectively mitigate the challenges posed by imbalanced datasets and erroneous labels, prevalent in real-world scenarios.
The positive impacts of this research are multifaceted. It enhances the accessibility and utility of machine learning in domains where data imbalance is a common challenge, such as healthcare, finance, and social media analytics. By improving class-balancedness in annotated sets, models trained on these datasets are less biased and more representative of real-world distributions, leading to fairer and more accurate outcomes. Additionally, this research contributes to reducing the time and cost associated with data annotation, which is particularly beneficial in fields where expert annotation is expensive or scarce.
However, if not carefully implemented, active learning strategies could inadvertently introduce new biases or amplify existing ones, particularly in scenarios where the initial data is severely imbalanced or contains deeply ingrained biases. Furthermore, the advanced nature of these techniques may widen the gap between organizations with access to state-of-the-art technology and those without, potentially exacerbating existing inequalities in technology deployment.
Appendix A Equivalent Objective
Lemma A.1.
The agnostic active learning reduction is equivalently finding the optimal separation threshold. Namely,
$\operatorname*{arg\,min}_{j}\mathcal{L}(h_{j})=\operatorname*{arg\,max}_{j}\left(\lvert\{y_{(i)}=1:1\leq i\leq j\}\rvert-\lvert\{y_{(i)}=2:1\leq i\leq j\}\rvert\right)$
Proof.
Recall the definitions: $h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}$ and $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$, we can expand the loss as follows
$\operatorname*{arg\,min}_{j}\mathcal{L}(h_{j})$ $=\operatorname*{arg\,min}_{j}\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$
$=\operatorname*{arg\,min}_{j}N-\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})=y_{(i)}\}$
$=\operatorname*{arg\,max}_{j}\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})=y_{(i)}\}$
$=\operatorname*{arg\,max}_{j}\left(\sum_{i=1}^{j}\mathbf{1}\{y_{(i)}=1\}\right)+\left(\sum_{i=j+1}^{N}\mathbf{1}\{y_{(i)}=2\}\right)$
$=\operatorname*{arg\,max}_{j}\left(\sum_{i=1}^{j}\mathbf{1}\{y_{(i)}=1\}\right)+\left(\sum_{i=j+1}^{N}\mathbf{1}\{y_{(i)}=2\}\right)-\left(\sum_{i=1}^{N}\mathbf{1}\{y_{(i)}=2\}\right)$
$=\operatorname*{arg\,max}_{j}\sum_{i=1}^{j}\left(\mathbf{1}\{y_{(i)}=1\}-\mathbf{1}\{y_{(i)}=2\}\right)$
Appendix B All Results
B.1 Noiseless Results under Imbalance
Figure 6: Imbalanced CIFAR-10, two classes.
Figure 7: Imbalanced CIFAR-10, three classes.
Figure 8: Imbalanced CIFAR-100, two classes.
Figure 9: Imbalanced CIFAR-100, three classes.
Figure 10: Imbalanced CIFAR-100, 10 classes.
Figure 11: Imbalanced SVHN, two classes.
Figure 12: Imbalanced SVHN, three classes.
Figure 13: PathMNIST, two classes.
(a) FMoW Balanced Pool Accuracy
(b) iWildcam Balanced Pool Accuracy
Figure 14: LabelBench results in the noiseless setting.
B.2 Label Noise Results under Imbalance
Figure 15: Imbalanced CIFAR-10, two classes, 10% label noise.
Figure 16: Imbalanced CIFAR-10, three classes, 10% label noise.
Figure 17: Imbalanced CIFAR-100, two classes, 10% label noise.
Figure 18: Imbalanced CIFAR-100, two classes, 15% label noise.
Figure 19: Imbalanced CIFAR-100, two classes, 20% label noise.
Figure 20: Imbalanced CIFAR-100, three classes, 10% label noise.
Figure 21: Imbalanced SVHN, two classes, 10% label noise.
Figure 22: Imbalanced SVHN, three classes, 10% label noise.
(a) FMoW Balanced Pool Accuracy
(b) iWildcam Balanced Pool Accuracy
Figure 23: LabelBench results in the 10% label noise setting.